OpenStudy (jodecci):

The hypotenuse of an isosceles triangles measures 10 inches long. What is the length of one leg of the triangle?

OpenStudy (aveline):

 In order to be an isosceles triangle with a hypotenuse, it must be a 45/45/90 triangle! \[x ^{2}+x ^{2}=10^{2}\] \[2x ^{2}=100\] \[x ^{2}=50\] \[x=\sqrt{50}\] \[x=\sqrt{25}\times \sqrt{2}\] \[x=5*\sqrt{2}\] \[x=5\sqrt{2}\]

OpenStudy (aveline):

@jodecci

OpenStudy (jodecci):

would 10 over 2 squar rooted be the same? @Aveline

OpenStudy (aveline):

This?\[\frac{ 10 }{ \sqrt{2} }\] or this?\[\sqrt{\frac{ 10 }{ 2 }}\]

OpenStudy (jodecci):

top one @Aveline

OpenStudy (aveline):

@jodecci It would be the same thing. However, your teacher may not want radicals in your denominator.

OpenStudy (jodecci):

okay thank you so much